# MV2Cyl: Reconstructing 3D Extrusion Cylinders from Multi-View Images

### v1

@@ 4.1 Results
MV2Cyl also outperforms Point2Cyl [58], which explicitly leverages ground truth, precise 3D geometric information as input. An advantage of taking multi-view images instead of point clouds is the images contain richer information such as edges and curves. Moreover, 2D neural networks have also shown superior performance compared to 3D processing networks leading to superior performance of MV2Cyl compared to Point2Cyl [58]. The comparison of the 2D segmentation network in MV2Cyl with the 3D segmentation network in Point2Cyl is detailed in the appendix A6.
Unlike the baselines that require accurate 3D surface information, MV2Cyl is able to directly handle 2D multi-view images as input by leveraging on learned 2D priors to extract 3D extrusion information. We prove that extracting extrusion surfaces and curves from 2D images allows us to leverage both the curve and surface representation for better extrusion cylinder reconstruction compared to baselines that even require clean 3D geometry as input.
@@ 4.2 Real Object Reconstruction
Finally, we also show the feasibility of MV2Cyl on a real world demo. We 3D-printed different sketch-extrude objects from test set of Fusion360 [63], and then captured multi-view images using an iPhone12 Mini. The images are then fed into COLMAP [52] to obtain camera poses. The background is then removed using off-the-shelf background removal. The images are then fed into our pipeline to extract the information from our 2D prior, then distilling this into 3D with our existence and attribute fields, and finally recovering extrusion cylinders and its parameters. More details and examples are found in the appendix A2.
@@ 5 Conclusion
We introduce MV2Cyl, a novel method for reconstructing a set of extrusion cylinders in 3D from multi-view images. Unlike previous approaches that rely only on raw 3D geometry as input, our framework demonstrates superior performance by fully leveraging the capabilities of 2D convolutional neural networks in image segmentation. Through the analysis of curve and surface representations, each has its advantages and shortcomings, we propose an integration of them that takes the best of both worlds, achieving optimal performance.
Limitations
While MV2Cyl achieves the state-of-the-art performance in reconstructing sketch-extrude CAD models, similar to prior work such as Point2Cyl [58], our framework is limited in predicting binary operations across primitives. We plan to further explore predicting binary operations using multi-view image inputs.
@@
Appendix A1 Ablation Study
Importance of Jointly Using Surface and Curve
Tab. A2 shows the importance of combining surface and curve reconstruction for CAD reconstruction. Fig. A4 shows qualitative results of the final CAD model outputs and also includes intermediate results, such as point clouds from existence fields.
@@
Figure A4: Ablation study. Each example illustrates the limitations of surface and curve representations. The surface-only model, whose instance label point cloud, start-end label point cloud, and output CAD model are shown in columns 2, 3, and 4, The table shows the final output of the surface-only model and the edge-only model along with their semantic fields. The surface-only model failed to reconstruct the proper start-end field and thus the sketch is invalid. The curve-only model was able to reconstruct the geometry in case (a) but predicted zero height for the thin cylinder in (b). Ours were able to make better estimations in both cases.
@@
Method E.A.$\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
@@
Appendix A2 Real-World Demo
@@
To reverse engineer a real-world object, we first capture videos of the target object using an iPhone12 Mini and convert these into multi-view images by extracting frames. We then preprocess the images by removing the background through luminance thresholding and setting it to black, and then we transform the images into grayscale. As the U-Net networks in our segmentation framework are trained on synthetic images rendered from ground truth meshes using Blender, there exists a domain gap that affects the segmentation performance on real images. To address this issue and improve segmentation results, we leverage a powerful 2D segmentation model called Segment Anything (SAM) [24]. By fine-tuning the pretrained SAM model using LoRA-based finetuning code [70] with the same training dataset used for training the U-Nets, we observed significant improvements in segmentation performance on real images despite the distributional shift. The finetuned SAM model benefits from its training on a significantly larger dataset, providing rich prior knowledge of 2D segmentation and ensuring robust performance. We acquire the camera’s extrinsic and intrinsic parameters using COLMAP [52, 53].
Figure A5: Example of curve segmentation prediction for the real images. The first image is the real image and it is pre-processed to the second image. SAM [24] segments the images into the existence map (the third), the instance segment (the fourth), and the start-end segment (the last image).
Fig. A5 displays an example of a real image alongside the predicted segmentation maps from the finetuned SAM model. The processed multi-view images and their corresponding camera poses are then input into the reconstruction pipeline, which focuses on learning the existence and attribute fields. Subsequently, CAD parameters are estimated based on this reconstruction. Tab. A3 presents results on final reconstructed CAD models in real captures, demonstrating successful output of CAD models from various real-world objects.
@@
Appendix A3 Comparison with SECAD-Net [30]
@@
Dataset Method E.A.$\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl. $\downarrow$ Fit. Glob.$\downarrow$
@@
Tab. A4 demonstrates that MV2Cyl outperforms SECAD-Net [30] across all metrics. SECAD-Net lacks supervision for segmenting parts in the way CAD designers would, leading to issues such as over-segmentation. This is evident in the images of Tab. A5. The second row shows that SECAD-Net divides two vertically attached cylinders into several patches along the cylinders, while MV2Cyl accurately represents them as two distinct cylinders.
@@
Appendix A4 Comparison with Curve + Point2Cyl [58]
@@
Method E.A.$\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
@@
Appendix A5 Implementation Details
We elaborate the implementation of the networks presented in the main paper. Sec. A5.1 contains the details of the training 2D segmentation frameworks (Sec. 3.2 in the main paper) and Sec. A5.2 includes the details of the 3D field learning from the 2D segments (Sec. 3.3 in the main paper).
A5.1 Details of 2D Segmentation Frameworks
@@
A5.2 Details of Integrating 2D Segments to 3D Fields
The opacity field $\sigma(\textbf{x})$ hyper-parameters are set to $\zeta=10$ and $\beta=0.8$ as in [67]. For the reconstruction loss of the existence field $\mathcal{F}$, we set $\lambda_{sparsity}=0.5$, $s=0.5$, $\eta_{\text{batch}}=0.1$, and $\eta_{\text{image}}=0.05$. We use a batch size of 8,192 rays for the training of both the existence field $\mathcal{F}$ and the attribute field $\mathcal{A}$. We trained all the fields for 1500 iterations on a single NVIDIA RTX 3090 GPU. The whole training time is approximately 5 minutes per shape. Note that the segmentation process and the learning of object existence and attributes are typically done as separate steps, rather than in a single, continuous process.
A5.3 Reverse Engineering Implementation Details
We provide additional details on our reverse engineering process, i.e. from our existence and attribute fields to the reconstruction of the extrusion cylinders, as discussed in Sec. 3.4 of the main paper. This process begins by sampling points from the existence fields, and then from the extracted point clouds deriving the corresponding parameters for each extrusion cylinder. We query the learned existence field value across a $400\times 400\times 400$ grid. Points on this grid with an opacity $\sigma$ value greater than a threshold $\tau\geq 0.99$ are selected and their corresponding instance segmentation and start-end(-barrel) segmentation labels are extracted.
The extrusion axis ($\mathbf{n}$), height ($\mathbf{h}$), and center ($\mathbf{c}$) are determined directly from the sampled point clouds as detailed in the main paper. The process to extract the 2D implicit sketch involves additional steps detailed as follows: After predicting the extrusion axis, the 2D sketch boundary points are created by projecting the base curve points along this axis. These 2D point clouds are then used to infer the 2D implicit sketch using IGR [15], which takes as input the projected 2D sketch point clouds and their normals to generate the corresponding signed distance fields. To extract the 2D point normals for IGR, we modify PGR [32] to estimate 2D normals from our projected 2D sketch points. We adapt the algorithm from PGR and modified 2D basis as follows:
@@
A5.4 Multi-View Images Preparation Details
@@
A5.5 NeuS2 [61] + Point2Cyl [58] Implementation Details
@@
Appendix A6 Additional Analysis
Comparison of Segmentation Performance Between 2D and 3D Networks
@@
Appendix A7 Additional Qualitative Results

### v2

@@ 4.1 Results
MV2Cyl also outperforms Point2Cyl [58], which takes precise 3D geometric information as input. An advantage of taking multi-view images instead of point clouds is the images contain richer information such as edges and curves. Moreover, 2D neural networks have also shown superior performance compared to 3D processing networks leading to superior performance of MV2Cyl compared to Point2Cyl [58]. The comparison of the 2D segmentation network in MV2Cyl with the 3D segmentation network in Point2Cyl are found in Appendix F.1.
Unlike the baselines that require accurate 3D surface information, MV2Cyl is able to directly handle 2D multi-view images as input by leveraging on learned 2D priors to extract 3D extrusion information. We show that extracting extrusion surfaces and curves from 2D images allows us to leverage both the curve and surface representation for better extrusion cylinder reconstruction compared to baselines that even require clean 3D geometry as input.
@@ 4.2 Real Object Reconstruction
We validate MV2Cyl on a real-world demo by 3D-printing various sketch-extrude objects from the Fusion360 test set, by capturing multi-view images with an iPhone 12 Mini and then obtaining camera poses via COLMAP. These captured images were processed through our pipeline to extract 2D information, converted to 3D using our density and semantic fields to recover extrusion cylinders and parameters. To address the domain gap between synthetic training images and real-world data, we applied preprocessing techniques, including grayscale conversion, background removal, and then fine-tuning a large vision model. Additional details are provided in Appendix B.
@@ 5 Conclusion
We introduce MV2Cyl, a novel method for reconstructing a set of extrusion cylinders in 3D from multi-view images. Unlike previous approaches that rely only on raw 3D geometry as input, our framework utilized 2D multi-view images and demonstrates superior performance by leveraging the capabilities of 2D convolutional neural networks in image segmentation. Through the analysis of curve and surface representations, each having its own advantages and shortcomings, we propose an integration of both that takes the best of both worlds, achieving optimal performance.
Limitations and Future Work
Our proposed method is not without its limitations. While MV2Cyl demonstrates state-of-the-art performance in reconstructing sketch-extrude CAD models, it shares a limitation with previous approaches, such as Point2Cyl [58], regarding the prediction of binary operations among primitives. MV2Cyl does not explicitly predict these binary operations, which we leave as future work. Nonetheless, we propose an initial straightforward method for recovering binary operations through exhaustive search. Details are provided in the appendix E.6. Additionally, MV2Cyl does not directly accommodate textured CAD models, as the 2D segmentation models are trained on a synthetic dataset derived from rendering untextured CAD models. By utilizing generalizable large 2D models, such as Segment Anything [24], we can preprocess images of textured CAD models for segmentation by extracting object masks and removing textures. We also leave this for future investigation.
MV2Cyl is also susceptible to occlusion since it relies on 2D image inputs. Specifically, if one side of an extrusion cylinder is completely hidden by others, our 2D segmentation model cannot detect the hidden side. In such cases, the extrusion cylinder cannot be reconstructed. In Fig. 6, the left shape represents the target CAD model, while the right one shows the reconstructed model by MV2Cyl. The target shape features an inset hexagonal cylinder with one end hidden by an outer cylinder, but MV2Cyl was unable to reconstruct the inset extrusion cylinder.
Figure 6: Example of a failure case.
@@
Acknowledgements
We thank Seungwoo, Juil, Jaihoon, and Yuseung for their valuable comments and suggestions in developing the idea and preparing an earlier version of the manuscript. This work was supported by the NRF grant (RS-2023-00209723), IITP grants (RS-2022-II220594, RS-2023-00227592, RS-2024-00399817), and KEIT grant (RS-2024-00423625), all funded by the Korean government (MSIT and MOTIE), as well as grants from the DRB-KAIST SketchTheFuture Research Center, NAVER-Intel Co-Lab, Hyundai NGV, KT, and Samsung Electronics. Mikaela also acknowledges the support from an Apple AI/ML PhD Fellowship.
@@ Acknowledgements
Appendix A Ablation Study
A.1 Importance of Jointly Using Surface and Curve
Tab. A2 shows the importance of combining surface and curve reconstruction for CAD reconstruction. Fig. A7 shows qualitative results of the final CAD model outputs and also includes intermediate results, such as point clouds from density fields.
@@ A.1 Importance of Jointly Using Surface and Curve
Figure A7: Ablation study. Each example illustrates the limitations of surface and curve representations. The surface-only model, whose instance label point cloud, start-end label point cloud, and output CAD model are shown in columns 2, 3, and 4, The table shows the final output of the surface-only model and the curve-only model along with their semantic fields. The surface-only model failed to reconstruct the proper start-end field and thus the sketch is invalid. The curve-only model was able to reconstruct the geometry in case (a) but predicted zero height for the thin cylinder in (b). Ours were able to make better estimations in both cases.
@@
Method E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
@@
A.2 Separate 2D Segmentation Modules for Curve and Surface
In MV2Cyl, the curve and surface segmentation modules are kept separate. To evaluate how segmentation performance changes when the 2D curve and surface segmentation networks are combined, we test a shared U-Net backbone that branches into two outputs for curve and surface segmentation. The results of using merged and separate U-Nets to extract 2D segmentation maps for our reverse engineering extrusion cylinders task are shown in Tab. A3. While a slight improvement in performance is observed on some metrics (Fit Cyl. and Fit Glob.), the results from both approaches are comparable.
Table A3: Quantitative comparison with the U-Nets of surface and curve segmentation using Fusion360 dataset.
Method E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
Combined U-Net 1.4425 0.0397 0.1543 0.0279 0.0200
Separate U-Nets (Ours) 1.3939 0.0385 0.1423 0.0284 0.0212
A.3 Ablation on the Number of Instances
We evaluate the impact of varying the number of instances on MV2Cyl’s performance using the Fusion360 dataset. Tab. A4 shows the results breakdown of the models in the Fusion360 test set across different numbers of extrusion instances K. Note that the number of samples for K = 7 and K = 8 is very small, and these statistics should therefore be interpreted sparingly.
Table A4: Performance over the changes of the number of instances using Fusion360 dataset.
Number of instances Number of samples E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
1 250 0.0000 0.0011 0.0615 0.0169 0.0169
2 494 1.4484 0.0325 0.1531 0.0277 0.0224
3 207 2.9046 0.0732 0.2092 0.0405 0.0261
4 93 1.8476 0.1092 0.2270 0.0466 0.0300
5 40 2.4250 0.0998 0.2065 0.0586 0.0294
6 15 9.0000 0.1301 0.1916 0.0601 0.0459
7 7 1.3265 0.0712 0.0620 0.0553 0.0364
8 2 16.2143 0.0130 0.0350 0.0250 0.0232
A.4 Ablation on the Number of the Input Images
In this section, we analyze the effect of varying the number of input images on MV2Cyl’s performance using the Fusion360 dataset. Tab. A5 shows that our approach is robust to changes in the number of input images, maintaining reasonable performance even with as few as 10 input views. Although performance slightly degrades as the number of input images decreases, MV2Cyl still produces effective reconstructions, demonstrating its adaptability to limited input data. As expected, the use of 50 input images yields the best overall results due to the richer information available for reconstruction.
Table A5: Performance over the changes of the number of the input images using Fusion360 dataset.
Number of images E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
50 (Ours) 1.3939 0.0385 0.1423 0.0284 0.0212
25 1.4631 0.0421 0.1561 0.0293 0.0218
15 2.1548 0.0489 0.1784 0.0371 0.0259
10 2.2585 0.0573 0.2112 0.0515 0.0364
A.5 Ablation on the Line Width of the Curve Segmentation Maps
In this section, we investigate the impact of varying the line width of curve segmentation maps on MV2Cyl’s performance. Tab. A6 shows that our method is relatively robust to changes in line width, with performance remaining consistent across different widths. The results indicate that using 5 pixels (our default setting) achieves a good balance between accuracy and consistency. While slight variations in error metrics are observed when using line widths of 2.5 or 7.5 pixels, these changes are minimal, suggesting that MV2Cyl can handle a range of line width settings without significant performance loss.
Table A6: Performance over the changes of the line width of the curve segmentation maps using Fusion360 dataset.
Line width E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
5 pixels (Ours) 1.3939 0.0385 0.1423 0.0284 0.0212
2.5 pixels 1.4886 0.0392 0.1482 0.0291 0.0216
7.5 pixels 1.4375 0.0410 0.1548 0.0307 0.0225
Appendix B Real-World Demo
@@
To reverse engineer a real-world object, we begin by capturing videos of the target object using an iPhone 12 Mini. These videos are then converted into multi-view images through frame extraction, which are subsequently input into the image segmentation models. The training dataset for our segmentation models is generated by rendering untextured models in Blender, producing RGB images with a grayscale appearance against a transparent background. To address the domain gap between the training data and real-world demo images, we preprocess the real images by converting them to grayscale and removing the background. While these preprocessing steps help align the datasets, a residual domain gap remains, which may still impact segmentation performance on real images. To address this challenge and enhance segmentation efficacy, we employ a robust 2D segmentation model known as Segment Anything (SAM)[24]. By fine-tuning the pretrained SAM model using LoRA-based finetuning code[71], with the same training dataset applied to the U-Nets, we observe substantial improvements in segmentation performance on real images, even in the presence of distributional shifts. The finetuned SAM model leverages extensive prior knowledge gained from training on a significantly larger dataset, ensuring robust performance in 2D segmentation tasks. We acquire the camera’s extrinsic and intrinsic parameters using COLMAP [52, 53].
Figure A8: Example of the inferred segmentation maps for a real image (top row) and a synthetic image (bottom row). The first image in the top row is the real image, which is then grayscaled (second image) and background-removed (third image). The fine-tuned SAM [24] model segments the processed image into the instance segment (fourth column) and the start-end segment (last column). When comparing the real and synthetic images from similar viewpoints, the segmentation outcomes are similarly effective.
Fig. A8 displays an example of a real image alongside the predicted segmentation maps from the finetuned SAM model. The processed multi-view images and their corresponding camera poses are then input into the reconstruction pipeline, which focuses on learning the density and semantic fields. Subsequently, CAD parameters are estimated based on this reconstruction. Tab. A7 presents results on final reconstructed CAD models in real captures, demonstrating successful output of CAD models from various real-world objects.
@@
We quantitatively evaluate MV2Cyl on real data found in Tab. A7. Our reconstruction achieves an average Chamfer distance of $2.11\times 10^{-3}$ from the ground truth CAD model. To compute this distance, we first sampled 8192-point point clouds from the reconstructions and then aligned the reconstructed model to the corresponding ground truth shape using ICP. After alignment, the desired metrics were computed. This process is illustrated in Fig. A9.
Figure A9: Processing steps for aligning the pose of the real demo output with the ground truth CAD point cloud.
Appendix C Comparison with SECAD-Net [30]
@@
Dataset Method E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl. $\downarrow$ Fit. Glob.$\downarrow$
@@
Tab. A8 demonstrates that MV2Cyl outperforms SECAD-Net [30] across all metrics. SECAD-Net lacks supervision for segmenting parts in the way CAD designers would, leading to issues such as over-segmentation. This is evident in the images of Tab. A9. For example, in the second row, SECAD-Net segments two vertically attached cylinders into several patches along their surfaces. In comparison, MV2Cyl more accurately represents these components as two distinct cylinders, showing its improved ability to align with CAD design conventions.
@@
Appendix D Comparison with Curve + Point2Cyl [58]
@@
Method E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl.$\downarrow$ Fit. Glob.$\downarrow$
@@
Appendix E Implementation Details
In this section, we elaborate the implementation of the networks presented in the main paper. Sec. E.1 contains the details of the training 2D segmentation frameworks (Sec. 3.2 in the main paper) and Sec. E.2 includes the details of the 3D field learning from the 2D segments (Sec. 3.3 in the main paper).
E.1 Details of 2D Segmentation Frameworks
@@
E.2 Details of Integrating 2D Segments to 3D Fields
The opacity field $\sigma(\textbf{x})$ hyper-parameters are set to $\zeta=10$ and $\beta=0.8$ as in [68]. For the reconstruction loss of the density field $\mathcal{F}$, we set $\lambda_{sparsity}=0.5$, $s=0.5$, $\eta_{\text{batch}}=0.1$, and $\eta_{\text{image}}=0.05$. We use a batch size of 8,192 rays for the training of both the density field $\mathcal{F}$ and the semantic field $\mathcal{A}$. We trained all the fields for 1500 iterations on a single NVIDIA RTX 3090 GPU. The whole training time is approximately 5 minutes per shape. Note that the segmentation process and the learning of object density and semantics are typically done as separate steps, rather than in a single, continuous process.
E.3 Reverse Engineering Implementation Details
We provide additional details on our reverse engineering process, i.e. from our density and semantic fields to the reconstruction of the extrusion cylinders, as discussed in Sec. 3.4 of the main paper. This process begins by sampling points from the density fields, and then from the extracted point clouds deriving the corresponding parameters for each extrusion cylinder. We query the learned density field value across a $400\times 400\times 400$ grid. Points on this grid with an opacity $\sigma$ value greater than a threshold $\tau\geq 0.99$ are selected and their corresponding instance segmentation and start-end(-barrel) segmentation labels are extracted. The extrusion axis ($\mathbf{n}$), height ($\mathbf{h}$), and center ($\mathbf{c}$) are determined directly from the sampled point clouds as detailed in the main paper.
The process to extract the 2D implicit sketch involves additional steps detailed as follows: After predicting the extrusion axis, the 2D sketch boundary points are created by projecting the base curve points along this axis. These 2D point clouds are then used to infer the 2D implicit sketch using IGR [15], which takes as input the projected 2D sketch point clouds and their normals to generate the corresponding signed distance fields. To extract the 2D point normals for IGR, we modify PGR [32] to estimate 2D normals from our projected 2D sketch points. We adapt the algorithm from PGR and modified 2D basis as follows:
@@ E.3 Reverse Engineering Implementation Details
E.4 Multi-View Images Preparation Details
@@
E.5 NeuS2 [62] + Point2Cyl [58] Implementation Details
@@
E.6 Binary Operation Prediction
We introduce a straightforward method for recovering binary operations. A naïve approach involves performing an exhaustive search across all possible combinations of primitives and operations, resulting in $2^{K}$ possibilities for a model with $K$ primitives. To identify the optimal combination that aligns closely with the observed input images, we measure the average per-pixel L2 distance between the rendered images of each combination’s resulting model and the observed multi-view images, selecting the combination with the lowest distance. We have implemented this approach to recover the binary operations, and examples of the results are illustrated in Fig. A13.
Figure A13: Example of the binary operation prediction of the reconstructed extrusion cylinders.
E.7 Implicit sketch to parametric curves
We also provide example showcases of converting an implicit sketch into a parametric representation. After generating the 2D implicit field for each sketch, we sample boundary points using the marching squares algorithm, which serve as knots for cubic Bézier curves. Using these points, we employ an off-the-shelf spline fitting module from the SciPy library [59] to determine the optimal B-spline representation based on the provided control and knot points. This process is illustrated in Fig. A14. It is important to note that the resulting sketch may not comprise the minimal set of parametric curves; for example, a single line segment may be represented by multiple segments.
Figure A14: Recovered parametric models with sketch represented as cubic spline.
Appendix F Additional Analysis
F.1 Comparison of Segmentation Performance Between 2D and 3D Networks
@@
F.2 Consistency Between Curve and Surface Segmentations
We compute a consistency metric between the curve and surface predictions. Specifically, we compute the percentage of models where the number of predicted curve instances and surface instances are identical. Across all the test shapes in the Fusion360 dataset, 98% of the models (1,095 shapes over 1,108 shapes) have the same number of curve and surface segments. Moreover, we also measure the consistency between the extracted curve and surface point clouds. Concretely, we measure the average one-way Chamfer distance between the extracted curve point cloud and surface point cloud ($0.081\times 10^{-3}$). We see that this very small value suggests that the extracted curve point clouds are quite consistent with the surface point clouds.
F.3 Chamfer Distance Evaluation for 3D Reconstruction
We quantitatively evaluate the 3D reconstruction quality of MV2Cyl by measuring the Chamfer distance against the ground truth CAD model. To generate the output CAD model, binary operations are determined through a simple exhaustive search approach. The results, shown in Tab. A12, demonstrate that our method outperforms the baseline, NeuS2+Point2Cyl. For a fair comparison, we compare with NeuS2+Point2Cyl instead of NeuS2, as our work focuses on the reverse engineering of extrusion cylinders, which is a more challenging task than 3D reconstruction without CAD structure. Reverse engineering extrusion cylinders enables shape editing and importing the model back into CAD software.
Table A12: Chamfer distances comparisons using Fusion360 and DeepCAD datasets. The chamfer distance values are multiplied by $10^{3}$.
Dataset Method Mean CD ($\downarrow$) Median CD ($\downarrow$)
Fusion360 NeuS2+Point2Cyl 105.18 51.88
MV2Cyl (Ours) 49.08 9.40
DeepCAD NeuS2+Point2Cyl 106.59 17.21
MV2Cyl (Ours) 18.21 1.43
Appendix G Additional Qualitative Results